OpenStudy (anonymous):

A cookie recipe requires 2.5 cups of sugar for every 10 cups of flour. If a baker makes enough cookies to use 16 cups of sugar, how many cups of flour would be used?

OpenStudy (mertsj):



OpenStudy (anonymous):

2.5/16=10/x cross multiply x=64 64 cups of flour.

OpenStudy (anonymous):

thank you both for your hard work